# Advancing African-Accented English Speech Recognition: Epistemic Uncertainty-Driven Data Selection for Generalizable ASR Models

### v6

@@ Abstract
Accents play a pivotal role in shaping human communication, enhancing our ability to convey and comprehend messages with clarity and cultural nuance. While there has been significant progress in Automatic Speech Recognition (ASR), African-accented English ASR has been understudied due to a lack of training datasets, which are often expensive to create and demand colossal human labor. Combining several active learning paradigms and the core-set approach, we propose a new multi-rounds adaptation process that uses epistemic uncertainty to automate the annotation process, significantly reducing the associated costs and human labor. This novel method streamlines data annotation and strategically selects data samples contributing most to model uncertainty, enhancing training efficiency. We define a new U-WER metric to track model adaptation to hard accents. We evaluate our approach across several domains, datasets, and high-performing speech models. Our results show that our approach leads to a 27% WER relative average improvement while requiring on average 45% less data than established baselines. Our approach also improves out-of-distribution generalization for very low-resource accents, demonstrating its viability for building generalizable ASR models in the context of accented African ASR. We open-source the code here.
@@ 1 Introduction
Automatic Speech Recognition (ASR) is an active research area that powers voice assistant systems (VASs) like Siri and Cortana, enhancing daily communication [27, 17, 46]. Despite this progress, no current VASs include African languages, which account for about 31% of the world languages, and their unique accents [16, 44]. This gap underscores the need for ASR systems that can handle the linguistic diversity and complexity of African languages, especially in crucial applications like healthcare. Due to the lack of representations of these languages and accents in training data, existing ASR systems often perform inadequately, even mispronouncing African names [36].
@@
To address these challenges, our work focuses on adapting pretrained speech models to better transcribe African-accented English, defined by unique intonations and pronunciations [7, 22]. We use epistemic uncertainty (EU) [25] to guide the adaptation process by identifying gaps in model knowledge and prioritizing data for the model to learn from next. This is particularly beneficial in scenarios where data annotation is costly or time-consuming, as often seen in the African context [5, 4, 6, 45, 12, 14, 13]. EU also improves robustness and encourages exploration to mitigate inductive bias from underrepresented accents. Common approaches to compute EU include Monte Carlo Dropout (MC-Dropout) [18] and Deep Ensembles [30], with the latter being more effective but computationally expensive. Due to resource constraints, we use MC-Dropout, which requires models to have dropout components during pretraining.
@@
To further enhance the efficiency and effectiveness of model adaptation, we employ Active Learning (AL) techniques. AL leverages epistemic uncertainty to select the most informative data points from an unlabeled dataset for labeling, thereby improving model performance with fewer training instances. Common types of AL include Deep Bayesian Active Learning (DBAL) [19, 23] and Adversarial Active Learning (AAL) [15]. AAL selects examples likely to be misclassified by the current model, refining it iteratively by challenging it with difficult cases to enhance robustness. The core-set approach (CSA) [42] is also related, as it selects a subset of training data to ensure a model trained on this subset performs comparably to one trained on the full dataset, addressing scalability and efficiency. A critical component of AL is the acquisition function (AF), which determines the most informative samples from an unlabeled dataset for labeling. Key AFs include uncertainty sampling (US) [31], Bayesian Active Learning by Disagreement (BALD) [19], and BatchBALD [26]. US targets data points with the highest model uncertainty. BALD maximizes the mutual information between model parameters and predictions. BatchBALD is an extension of BALD that selects multiple samples simultaneously but may choose redundant points. US is the least computationally expensive, making it ideal for efficient data labeling.
@@
In this work, we leverage and combine DBAL, AAL, US, and CSA in the following way (in order): First, we integrate the CSA by leveraging smaller training subsets ($\sim 45\%$ smaller than the full available training sets). Second, we use DBAL with MC-Dropout, to apply dropout during training and inference to estimate Bayesian posterior distribution. This allows us to practically and efficiently estimate EU in the models used [19] (see section 3.2 for more details). Third, we use the estimated EU and integrate the idea of AAL by using the US acquisition function.
@@
We evaluate our approach across several domains (general, clinical, general+clinical aka both), several datasets (AfriSpeech-200 [37], SautiDB [2], MedicalSpeech, CommonVoices English Accented Dataset [3]), and several high-performing speech models (Wav2Vec2-XLSR-53 [10], HuBERT-Large [24], WavLM-Large [8], and NVIDIA Conformer-CTC Large (en-US) [20]). Our results show a 27% Word Error Rate (WER) relative average improvement while requiring on average 45% less data than established baselines. We also adapt the standard WER to create a new metric called Uncertainty WER (U-WER) to track model adaptation to African accents.
@@
The impact of our approach is substantial. It develops more robust, generalizable, and cost-efficient African-accented English ASR models and reduces dependency on large labeled datasets, enabling deployment in various real-world scenarios. Our results show improved generalization for out-of-distribution (OOD) cases, especially for accents with minimal resources, addressing specific challenges in African-accented ASR. Additionally, by focusing on equitable representation in ASR training, our methodology promotes fairness in AI, ensuring technology serves users across diverse linguistic backgrounds without bias [41, 34, 32]. Our contributions are listed as follows:
@@
  • we evaluate our approach across several speech domains (clinical, general, both), and African-accented speech datasets (AfriSpeech-200 [37], SautiDB [2], MedicalSpeech and CommonVoices English Accented Dataset [3]), while providing domain and accent-specific analyses,
@@ 2.1 Challenges for African-accented ASR
State-of-the-art (SOTA) ASR technologies, powered by deep learning and neural network architectures like transformers, achieve high accuracy with Standard American English and major European languages. However, they often fail with African accents due to high variability in pronunciation and lack of quality speech data [29, 11]. This results in racial bias, poor performance, and potential social exclusion as speakers might alter their speech to be understood [29, 28, 9, 33]. Enhancing ASR for African languages is crucial for equitable voice recognition, especially in healthcare, education, and customer service. Solutions should focus on diversifying training datasets and developing robust modeling techniques tailored to the unique characteristics of these languages.
@@ 2.2 Active Learning
AL aims to reduce the number of labeled training examples by automatically processing the unlabeled examples and selecting the most informative ones concerning a given cost function for a human to label. It is particularly effective when labeled data is scarce or expensive, optimizing the learning process by focusing on samples that most improve the model performance and generalization [43, 19]. Several works have demonstrated its effectiveness and efficiency. An AL setup involves an unlabeled dataset $\mathcal{D}_{\mathrm{pool}}=\{\mathbf{x}_{i}\}_{i=1}^{n_{\mathrm{pool}}}$, a labeled training set $\mathcal{D}_{\mathrm{train}}=\{\mathbf{x}_{i},\mathrm{y}_{i}\}_{i=1}^{n_{\mathrm{train}}}$, and a predictive model with likelihood $p_{w}(y|x)$ parameterized by $w\sim p(W|\mathcal{D}_{\mathrm{train}})$ ($W$ are the parameters of the model). The setup assumes the presence of an oracle to provide predictions $\mathrm{y}$ for all $x_{i}\in\mathcal{D}_{\mathrm{pool}}$. After training, a batch of data $\{\mathbf{x}^{*}_{i}\}_{i=1}^{b}$ is selected from $\mathcal{D}_{\mathrm{pool}}$ based on its EU.
In [21], AL was applied to a toy dataset of How May I Help You recordings. Confidence scores were estimated for each word and used to compute the overall confidence score for the audio sample. This approach achieved competitive results using 27% less data compared to the baseline. In [40], the authors estimated confidence scores for each utterance using an online algorithm with the lattice output of a speech recognizer. The utterance scores were filtered through an informativeness function to select an optimal subset of training samples, reducing the labeled data needed for a given WER by over 60%. [35] experimented with AL for accent adaptation in speech recognition. They adapted a source recognizer to the target accent by selecting a small, matched subset of utterances from a large, untranscribed, multi-accented corpus for human transcription. They used a cross-entropy-based relevance measure alongside uncertainty-based sampling. However, their experiments on Arabic and English accents showed worse performance compared to baselines, while using more hours of recordings.
@@ 3.1 Datasets
We used the AfriSpeech-200 dataset [37], a 200-hour African-accented English speech corpus for clinical and general ASR. This dataset includes 120+ African accents from five language families: Afro-Asiatic, Indo-European, Khoe-Kwadi (Hainum), Niger-Congo, and Nilo-Saharan, representing African regional diversity. It was crowd-sourced from over 2000 African speakers from 13 anglophone countries in sub-Saharan Africa and the US (see Table 1).
To demonstrate the dataset-agnostic nature of our approach, we also explored three additional datasets: (1) SautiDB [1], Nigerian accent recordings with 919 audio samples at a 48kHz sampling rate, totaling 59 minutes; (2) MedicalSpeech [^1] https://www.kaggle.com/datasets/paultimothymooney/medical-speech-transcription-and-intent, containing 6,661 audio utterances of common medical symptoms, totaling 8 hours; and (3) CommonVoices English Accented Dataset, a subset of English Common Voice (version 10) [3], excluding western accents to focus on low-resource settings.
@@ 3.2 Methodology
2:we define a list variable called wer_list and a dictionary variable called wer_target_dict respectively tracking all pairwise WERs and the average pairwise WER of each target prediction
@@
where $\mathbf{M}_{t}$ is a binary mask matrix sampled independently for each pass. EU(($x|g,T$)) can then be estimated from the $T$ stochastic forward passes as follows:
@@
The use of MC-Dropout requires models to have dropout components during training. This excludes some models like Whisper [38], which we still finetuned and evaluated as a baseline. We use four state-of-the-art pretrained models: Wav2Vec2-XLSR-53, HuBERT-Large, WavLM-Large, and NVIDIA Conformer-CTC Large (en-US), referred to as Wav2Vec, Hubert, WavLM, and Nemo, respectively.
@@ 3.3 Experimental Design
For our experiments, we used 6 RTX8000 GPUs and 4 A100 GPUs. Training and evaluation were conducted over a month. Our models have approximately 311 million trainable parameters. Each audio sample was normalized and processed at a 16kHz sample rate. We used default parameters from the HuggingFace library for each pretrained model.
Table 3: We used Wav2Vec to conduct initial experiments across domains and strategies to identify the best selection strategy. Models marked with ** are used to demonstrate that our algorithm is model agnostic, utilizing the EU-Most selection strategy, which has been proven the most effective. Our AL experiments also use this strategy. Wav2Vec, using the random strategy, scored 0.1111, 0.3571, and 0.1666 for the general, clinical, and both domains, respectively. We omit random results to enhance readability.
@@
Table 4: WER Evaluation Results on External Datasets, with $\alpha\in[0.60,0.65]$ as described in Section 3.1 and on Figure 1. We see an improvement for WER using our approach in all datasets, showing that our algorithm is dataset-agnostic.
@@
CommonVoices English Accented Dataset (v10.0) [3] 26614 62100 10350 232 5 0.50 0.22
@@ 4 Results and Discussion
To assess the performance improvement for each domain, we compute the relative average improvement $\text{RIA}_{wer,d}=\left(\frac{b_{wer}^{d}-s_{wer}^{d}}{b_{wer}^{d}}\right)\times 100\%$ where $b_{wer}^{d}$ and $s_{wer}^{d}$ are the average WER respectively of the baseline, and the best selection strategy, in a domain $d\in\{general,clinical,both\}$. A higher percentage reflects a higher improvement of our approach.
Table 3 shows the results of our experiments, indicating that our uncertainty-based selection approach significantly outperforms the baselines across all models, domains, and datasets: general (27.00%), clinical (15.51%), and both (26.56%). Our approach also surpasses Whisper-Medium [37, 39], demonstrating the importance of epistemic uncertainty in ASR for low-resource languages. The EU-Most selection strategy proves to be the most effective across all domains due to the model exposure to highly uncertain samples, enhancing robustness and performance. However, performance disparities between general and clinical domains are noted, likely due to clinical samples complexity. These findings confirm EU-Most as the superior selection strategy, as detailed in the results and illustrated in Figures 2, 3, and 4. This answers question 2.
To identify the best learning signals within a diverse dataset characterized by various accents, speaker traits, genders, and ages, we analyzed the top-k uncertain accents using the EU-Most selection strategy. Our findings, illustrated in Figures 2, 3, and 4, show that the top-10 accents (most represented in recording hours) remained consistently challenging across all rounds of analysis (refer to Figures 2, 3, 4 and Tables 6, 7, and 8). These accents, characterized by high linguistic richness and variability, aid in model learning and enhance performance over time. We positively answer questions 1 and 3, confirming that the model adapts effectively to the beneficial accents from all domains. This demonstrates that the model adapts qualitatively and quantitatively well to the beneficial accents and benefits from all domains. Figures 2 (b), 3 (b), and 4 (b) also affirm positive outcomes for question 4, showing consistent improvement or stable performance on low-resource accents. This highlights the relevance of our approach in addressing the challenges associated with the low-resource availability typical of many African accents and languages.
To demonstrate the agnostic aspect of our approach, we evaluated it with three additional pretrained models (Hubert, WavLM, and Nemo) and three datasets containing accented speech in general and clinical domains, using only the EU-Most selection strategy. The results, shown in Tables 3 and 4, indicate that our uncertainty-based adaptation approach consistently outperforms baselines. This confirms that our approach applies to any model architecture and dataset, and allows us to answer positively question 5.

### v7

@@ Abstract
Accents play a pivotal role in shaping human communication, enhancing our ability to convey and comprehend messages with clarity and cultural nuance. While there has been significant progress in Automatic Speech Recognition (ASR), African-accented English ASR has been understudied due to a lack of training datasets, which are often expensive to create and demand colossal human labor. By combining several active learning paradigms and the core-set approach, we propose a new multi-round adaptation process that utilizes epistemic uncertainty to automate annotation, thereby significantly reducing associated costs and human labor. This novel method streamlines data annotation and strategically selects data samples that contribute most to model uncertainty, thereby enhancing training efficiency. We define a new U-WER metric to track model adaptation to hard accents. We evaluate our approach across several domains, datasets, and high-performing speech models. Our results show that our approach leads to a 27% WER relative average improvement while requiring, on average, 45% less data than established baselines. Our approach also improves out-of-distribution generalization for very low-resource accents, demonstrating its viability for building generalizable ASR models in the context of accented African ASR. We open-source the code here.
Advancing African-Accented English Speech Recognition: Epistemic Uncertainty-Driven Data Selection for Generalizable ASR Models
Bonaventure F. P. Dossou McGill University, Mila Quebec AI Institute bonaventure.dossou@mila.quebec
@@ 1 Introduction
Automatic Speech Recognition (ASR) is an active research area that powers voice assistant systems (VASs) like Siri and Cortana, enhancing daily communication Kodish-Wachs et al. (2018); Finley et al. (2018); Zapata and Kirkedal (2015). Despite this progress, no current VASs include African languages, which account for about 31% of the world languages, and their unique accents (Eberhard et al. (2019); Tsvetkov (2017)). This gap highlights the need for ASR systems that can effectively handle the linguistic diversity and complexity of African languages, particularly in critical applications such as healthcare. Due to the lack of representations of these languages and accents in training data, existing ASR systems often perform inadequately, even mispronouncing African names (Olatunji et al. (2023a)).
@@
To address these challenges, our work focuses on adapting pre-trained speech models to transcribe African-accented English more accurately, characterized by unique intonations and pronunciations Benzeghiba et al. (2007); Hinsvark et al. (2021). We use epistemic uncertainty (EU) Kendall and Gal (2017) to guide the adaptation process by identifying gaps in model knowledge and prioritizing data for the model to learn from next. This is particularly beneficial in scenarios where data annotation is costly or time-consuming, as often seen in the African context Badenhorst and De Wet (2019, 2017); Barnard et al. (2009); Yemmene and Besacier (2019); DiChristofano et al. (2022); Dossou et al. (2022); Dossou and Emezue (2021). EU also improves robustness and encourages exploration to mitigate inductive bias from underrepresented accents. Common approaches to compute EU include Monte Carlo Dropout (MC-Dropout) Gal and Ghahramani (2016) and Deep Ensembles Lakshminarayanan et al. (2017), with the latter being more effective but computationally expensive. Due to resource constraints, we utilize MC-Dropout, which necessitates that models incorporate dropout components during pretraining.
@@
We employ Active Learning (AL) techniques further to enhance the efficiency and effectiveness of model adaptation. AL leverages epistemic uncertainty to select the most informative data points from an unlabeled dataset for labeling, thereby improving model performance with fewer training instances. Common types of AL include Deep Bayesian Active Learning (DBAL) Gal et al. (2017); Houlsby et al. (2011) and Adversarial Active Learning (AAL) Ducoffe and Precioso (2018). AAL selects examples likely to be misclassified by the current model, refining it iteratively by challenging it with complex cases to enhance robustness. The core-set approach (CSA) Sener and Savarese (2017) is also related, as it selects a subset of the training data to ensure that a model trained on this subset performs comparably to one trained on the entire dataset, thereby addressing scalability and efficiency. A critical component of AL is the acquisition function (AF), which determines the most informative samples from an unlabeled dataset for labeling. Key AFs include uncertainty sampling (US) Liu and Li (2023), Bayesian Active Learning by Disagreement (BALD) Gal et al. (2017), and BatchBALD Kirsch et al. (2019). US targets data points with the highest model uncertainty. BALD maximizes the mutual information between model parameters and predictions. BatchBALD is an extension of BALD that selects multiple samples simultaneously but may choose redundant points. US is the least computationally expensive, making it ideal for efficient data labeling.
@@
In this work, we leverage and combine DBAL, AAL, US, and CSA in the following way (in order): First, we integrate the CSA by leveraging smaller training subsets ($\sim 45\%$ smaller than the entire available training sets). Second, we utilize DBAL with MC-Dropout to apply dropout during both training and inference, thereby estimating the Bayesian posterior distribution. This allows us to practically and efficiently estimate EU in the models used Gal et al. (2017) (see section 3.2 for more details). Third, we use the estimated EU and integrate the idea of AAL using the US acquisition function.
@@
We evaluate our approach across several domains (general, clinical, general+clinical aka both), several datasets (AfriSpeech-200 Olatunji et al. (2023b)), SautiDB Afonja et al. (2021b), MedicalSpeech, CommonVoices English Accented Dataset Ardila et al. (2019), and several high-performing speech models (Wav2Vec2-XLSR-53 Conneau et al. (2020), HuBERT-Large Hsu et al. (2021), WavLM-Large Chen et al. (2022), and NVIDIA Conformer-CTC Large (en-US) Gulati et al. (2020). Our results show a 27% Word Error Rate (WER) relative average improvement while requiring 45% less data than established baselines. We also adapt the standard WER to create an Uncertainty WER (U-WER) metric to track model adaptation to African accents.
@@
The impact of our approach is substantial. It develops more robust, generalizable, and cost-efficient African-accented English ASR models, reducing dependency on large labeled datasets and enabling deployment in various real-world scenarios. Our results demonstrate improved generalization for out-of-distribution (OOD) cases, particularly for accents with limited resources, addressing specific challenges in African-accented automatic speech recognition (ASR). Additionally, by focusing on equitable representation in ASR training, our methodology promotes fairness in AI, ensuring technology serves users across diverse linguistic backgrounds without bias Selbst et al. (2019); Mitchell et al. (2019); Mehrabi et al. (2021). Our contributions are listed as follows:
@@
  • we evaluate our approach across several speech domains (clinical, general, both), and African-accented speech datasets AfriSpeech-200 Olatunji et al. (2023b), SautiDB Afonja et al. (2021b), MedicalSpeech, and CommonVoices English Accented Dataset Ardila et al. (2019), while providing domain and accent-specific analyses,
@@ 2.1 Challenges for African-accented ASR
State-of-the-art (SOTA) ASR technologies, powered by deep learning and neural network architectures like transformers, achieve high accuracy with Standard American English and major European languages. However, they often fail with African accents due to high variability in pronunciation and lack of quality speech data Koenecke et al. (2020); Das et al. (2021). This results in racial bias, poor performance, and potential social exclusion as speakers might alter their speech to be understood Koenecke et al. (2020); Koenecke (2021); Chiu et al. (2018); Mengesha et al. (2021). Enhancing Automatic Speech Recognition (ASR) for African languages is crucial for achieving equitable voice recognition, particularly in healthcare, education, and customer service. Solutions should focus on diversifying training datasets and developing robust modeling techniques tailored to the unique characteristics of these languages.
@@ 2.2 Active Learning
AL aims to reduce the number of labeled training examples by automatically processing unlabeled examples and selecting the most informative ones, considering a given cost function, for a human to label. It is particularly effective when labeled data is scarce or expensive, optimizing the learning process by focusing on samples that most improve the model performance and generalization Settles (2009); Gal et al. (2017). Several works have demonstrated its effectiveness and efficiency. An AL setup involves an unlabeled dataset $\mathcal{D}_{\mathrm{pool}}=\{\mathbf{x}_{i}\}_{i=1}^{n_{\mathrm{pool}}}$, a labeled training set $\mathcal{D}_{\mathrm{train}}=\{\mathbf{x}_{i},\mathrm{y}_{i}\}_{i=1}^{n_{\mathrm{train}}}$, and a predictive model with likelihood $p_{w}(y|x)$ parameterized by $w\sim p(W|\mathcal{D}_{\mathrm{train}})$ ($W$ are the parameters of the model). The setup assumes the presence of an oracle to provide predictions $\mathrm{y}$ for all $x_{i}\in\mathcal{D}_{\mathrm{pool}}$. After training, a batch of data $\{\mathbf{x}^{*}_{i}\}_{i=1}^{b}$ is selected from $\mathcal{D}_{\mathrm{pool}}$ based on its EU.
In Hakkani-Tür et al. (2002), AL was applied to a toy dataset of How May I Help You recordings. Confidence scores were estimated for each word and used to compute the overall confidence score for the audio sample. This approach achieved competitive results using 27% less data compared to the baseline. In Riccardi and Hakkani-Tur (2005), the authors estimated confidence scores for each utterance using an online algorithm with the lattice output of a speech recognizer. The utterance scores were filtered through an informativeness function to select an optimal subset of training samples, reducing the labeled data needed for a given WER by over 60%. Nallasamy et al. (2012) experimented with AL for accent adaptation in speech recognition. They adapted a source recognizer to the target accent by selecting a small, matched subset of utterances from a large, untranscribed, multi-accented corpus for human transcription. They employed a cross-entropy-based relevance measure in conjunction with uncertainty-based sampling. However, their experiments on Arabic and English accents showed worse performance compared to baselines while using more hours of recordings.
@@ 3.1 Datasets
We used the AfriSpeech-200 dataset Olatunji et al. (2023b), a 200-hour African-accented English speech corpus for clinical and general ASR. This dataset comprises over 120 African accents from five language families: Afro-Asiatic, Indo-European, Khoe-Kwadi (Hainum), Niger-Congo, and Nilo-Saharan, representing the diversity of African regional languages. It was crowd-sourced from over 2000 African speakers from 13 anglophone countries in sub-Saharan Africa and the US (see Table 1).
To demonstrate the dataset-agnostic nature of our approach, we also explored three additional datasets: (1) SautiDB Afonja et al. (2021a), Nigerian accent recordings with 919 audio samples at a 48kHz sampling rate, totaling 59 minutes; (2) MedicalSpeech [^1] https://www.kaggle.com/datasets/paultimothymooney/medical-speech-transcription-and-intent, containing 6,661 audio utterances of common medical symptoms, totaling 8 hours; and (3) CommonVoices English Accented Dataset, a subset of English Common Voice (version 10) Ardila et al. (2019), excluding western accents to focus on low-resource settings.
@@ 3.2 Methodology
2:we define a list variable called wer_list and a dictionary variable called wer_target_dict, respectively tracking all pairwise WERs and the average pairwise WER of each target prediction
@@
where $\mathbf{M}_{t}$ is a binary mask matrix sampled independently for each pass. EU($x|g,T$) can then be estimated from the $T$ stochastic forward passes as follows:
@@
The use of MC-Dropout requires models to have dropout components during training. This exclusion applies to some models, such as Whisper Radford et al. (2022), which we still fine-tuned and evaluated as a baseline. We utilize four state-of-the-art pre-trained models: Wav2Vec2-XLSR-53, HuBERT-Large, WavLM-Large, and NVIDIA Conformer-CTC Large (en-US), referred to as Wav2Vec, HuBERT, WavLM, and Nemo, respectively.
@@ 3.3 Experimental Design
For our experiments, we utilized six RTX 8000 GPUs and four A100 GPUs. Training and evaluation were conducted over a period of one month. Our models have approximately 311 million trainable parameters. Each audio sample was normalized and processed at a 16 kHz sample rate. We used default parameters from the HuggingFace library for each pretrained model.
Table 3: We utilized Wav2Vec to conduct initial experiments across various domains and strategies, aiming to identify the optimal selection strategy. Models marked with ** are used to demonstrate that our algorithm is model agnostic, utilizing the EU-Most selection strategy, which has been proven to be the most effective. Our AL experiments also use this strategy. Wav2Vec, using the random strategy, scored 0.1111, 0.3571, and 0.1666 for the general, clinical, and both domains, respectively. We omit random results to enhance readability.
@@
Table 4: WER Evaluation Results on External Datasets, with $\alpha\in[0.60,0.65]$ as described in Section 3.1 and on Figure 1. We observe an improvement in WER using our approach across all datasets, indicating that our algorithm is dataset-agnostic.
@@
CommonVoices English Accented Dataset (v10.0) Ardila et al. (2019) 26614 62100 10350 232 5 0.50 0.22
@@ 4 Results and Discussion
To assess the performance improvement for each domain, we compute the relative average improvement
$\text{RIA}_{wer,d}=\left(\frac{b_{wer}^{d}-s_{wer}^{d}}{b_{wer}^{d}}\right)\times 100\%$
where $b_{wer}^{d}$ and $s_{wer}^{d}$ are the average WER respectively of the baseline, and the best selection strategy, in a domain $d\in\{general,clinical,both\}$. A higher percentage reflects a higher improvement in our approach.
Table 3 shows the results of our experiments, indicating that our uncertainty-based selection approach significantly outperforms the baselines across all models, domains, and datasets: general (27.00%), clinical (15.51%), and both (26.56%). Our approach also surpasses Whisper-Medium (Olatunji et al. (2023b); Radford et al. (2023)), demonstrating the importance of epistemic uncertainty in ASR for low-resource languages. The EU-Most selection strategy proves to be the most effective across all domains due to the model’s exposure to highly uncertain samples, enhancing robustness and performance. However, performance disparities between the general and clinical domains are noted, likely due to the complexity of the clinical sample. These findings confirm EU-Most as the superior selection strategy, as detailed in the results and illustrated in Figures 2, 3, and 4. This answers question 2.
To identify the best learning signals within a diverse dataset characterized by various accents, speaker traits, genders, and ages, we analyzed the top-k uncertain accents using the EU-Most selection strategy. Our findings, illustrated in Figures 2, 3, and 4, show that the top-10 accents (most represented in recording hours) remained consistently challenging across all rounds of analysis (refer to Figures 2, 3, 4 and Tables 6, 7, and 8). These accents, characterized by high linguistic richness and variability, facilitate model learning and improve performance over time. We positively answer questions 1 and 3, confirming that the model adapts effectively to the beneficial accents from all domains. This demonstrates that the model adapts qualitatively and quantitatively well to the beneficial accents and benefits from all domains. Figures 2 (b), 3 (b), and 4 (b) also affirm positive outcomes for question 4, showing consistent improvement or stable performance on low-resource accents. This highlights the relevance of our approach in addressing the challenges associated with the limited resource availability typical of many African languages and dialects.
To demonstrate the agnostic aspect of our approach, we evaluated it using three additional pre-trained models (Hubert, WavLM, and Nemo) and three datasets containing accented speech in general and clinical domains, employing only the EU-Most selection strategy. The results, shown in Tables 3 and 4, indicate that our uncertainty-based adaptation approach consistently outperforms baselines. This confirms that our approach applies to any model architecture and dataset, allowing us to answer question 5 positively.